# Tunable magnetism of Boron Imidazolate-based Metal-Organic Frameworks

### v2

@@ Abstract
Magnetic metal-organic frameworks (MMOFs), where magnetic metal nodes are connected into a crystal structure by organic linkers, have a potential to host exotic magnetic states. We present a study of bulk magnetic properties of four metal-organic frameworks with the same boron imidazolate linkers, Cu-BIF, Co-BIF, Ni-BIF, and newly synthesized Zn-BIF, displaying a variety of lattice structures and nontrivial magnetic behaviors. While non-magnetic Zn-BIF provides an offset of magnetic response, magnetic susceptibility measurements of the other three magnetic materials demonstrate the presence of weak magnetic interactions in these MOFs, which differ between materials by sign and size. Cu-BIF, where magnetic nodes are connected into octahedral cages, shows simple paramagnetic behavior. Triangular lattice Co-BIF shows antiferromagnetic interactions on the order of 1 K, and a spin-crossover-like effect in magnetic susceptibility due to thermal depopulation of excited crystal electric field levels. Magnetic properties of Ni-BIF suggest sizable ferromagnetic interactions. Using DC/AC susceptibility and variable-field DC magnetization, we detect cluster spin-glass behavior and discuss possible microscopic origins of this behavior. This work demonstrates the variety of magnetic properties that are possible with a single organic ligand, and establishes the low energy scale of magnetic interactions through superexchange in boron imidazolate frameworks.
@@ I Introduction
Frustrated magnetism is one of the hot topics of current condensed matter research. While one of the ultimate goals is to synthesize and characterize spin liquid materials, where highly interacting spins are not ordered down to low temperatures Balents2010 ; Broholm2020 , many exotic magnetic states and new materials are found as a part of this research. One underexplored source of new magnetic materials are magnetic metal-organic frameworks (MMOFs), where magnetic metal nodes are connected into crystals by organic linkers Thorarinsdottir2020 . Due to the versatility of organic chemistry and a relatively easy way to substitute metallic nodes without changing the structure or achieving different structures with the same magnetic nodes and linkers, these materials have the potential to be highly tunable magnets. There are a number of theoretical proposals on the realization of exotic magnetism in these materials Jacko2015 ; Yamada2017 ; yamada2017crystalline ; Jiang2021 ; zhang2017theoretical ; kambe2014redox , as well as a fair amount of MMOFs synthesized Thorarinsdottir2020 , but systematic studies of magnetic properties are rare.
In this manuscript, we present systematic magnetic studies of four boron imidazolate frameworks (BIFs), a unique subclass of MOFs that bear resemblances to zeolites and zeolitic imidazolate frameworks Zhang2016a ; Isaeva2020a , pictured in Fig. 1. Controlling the synthetic conditions as well as selecting the identity of the metal ion precursor can lead to a variety of solid-state structures. In this study, we selected 4 BIFs that are supported by the same boron trisimidazolate linker, but with different metal ions. Co-BIF and Ni-BIF are isostructural and form layered structures of 2D triangular lattices of MN6 octahedra (M = Co (J=1/2), Ni (S=1) ) with $D_{3d}$ symmetry connected by boron imidazolate molecules. Cu-BIF forms a M6L8 (M = metal ion, L = ligand) cage structure, with 6 Cu2+ ions (S=1/2) in an octahedral arrangement around a central pore. Finally, Zn-BIF forms a different cage structure (M4L4), with 4 Zn2+ (S=0) ions on opposite corners of a cube so as to form tetrahedra of Zn. These structures allow us to explore the effects of metal node substitution and MOF crystal structure on the magnetic properties of MMOFs.
@@ III.4 Ni-BIF
Figure 8: Top: in-phase component of AC susceptibility of Ni-BIF. $\chi^{\prime}$ data have been normalized to high-temperature $\chi^{\prime}$ at 999 Hz; raw data and scaling factors are presented in the SISI . Bottom left: Vogel-Fulcher fit to $T_{f}$ obtained from $\chi^{\prime}$. Bottom right: power law fit to $T_{f}$ obtained from $\chi^{\prime}$.
To further understand the temperature-dependent spin relaxation dynamics of Ni-BIF we performed ZFC AC susceptibility measurements on the same sample. In-phase susceptibility $\chi^{\prime}$ is shown in Fig. 8 for 5 different frequencies with zero applied DC field. The $\chi^{\prime}$ data have been normalized to the high-temperature susceptibility at 999 Hz in order to clearly show the evolution of the freezing temperature $T_{f}$. Raw data are presented in Fig. 7b of the SI SI . $\chi^{\prime}$ displays a broad, complex peak shape consistent with a broad crossover in DC magnetic susceptibility (Fig. 5). The peak in $\chi^{\prime}$ shifts to higher temperatures as frequency is increased. The out-of-phase component $\chi^{\prime\prime}$, shown in Fig. 7c of the SI SI , highlights the complex peak shape. $T_{f}$ is approximated from the maximal point on the high-temperature side of the broad peak, as shown in Fig. 8. $T_{f}$ changes from 11.7 K to 12.5 K between 1 Hz and 999 Hz; thus the Mydosh parameter, $\Delta T_{f}/(T_{f}\log_{10}(\Delta\nu))$, or the fractional change in $T_{f}$ per decade of frequency, is found to be 0.02. This parameter can be used to distinguish between spin-glass and superparamagnetic states, as in conventional spin-glasses it is found to be $\lesssim$ 0.01 while in superparamagnets it is $\gtrsim$ 0.1 Mydosh_2015 ; Bag2018 . An intermediate value such as that observed for Ni-BIF is attributed to cluster spin-glass behavior Bag2018 ; Banerjee2022_CSG .
An estimation of characteristic relaxation times allows to distinguish between spin glass and cluster spin glass Bag2018 ; Banerjee2022_CSG . Relaxation time can be estimated using the Vogel-Fulcher model (see Fig. 8) which adapts the noninteracting Néel-Arrhenius law with a phenomenological parameter $T_{0}$ accounting for intercluster interactions Souletie1985 :
@@
where $\tau^{*}$ is a characteristic relaxation time similar to $\tau_{0}$, $T_{g}$ is the spin-glass temperature at zero frequency, and the power $zv^{\prime}$, which is treated as a single value for the fit, is the dynamical critical exponent for the relaxation time near the crossover. The fit in Fig. 8 suggests $\log_{10}(\tau^{*})=-6.8(7)$ ($\tau^{*}\sim 4.1(3.8)*10^{-7}$ s), $T_{g}$ = 11.62(4) K, and $zv^{\prime}$ = 2.8(6) K. The suggested ranges of the characteristic relaxation time and the dynamical critical exponent overlap with the ranges of $10^{-7}-10^{-10}$ s and 3-10 typically observed in cluster spin-glass systems Banerjee2022_CSG .
@@ IV Discussion
A combination of magnetic frustration, competing anti- and ferromagnetic interactions, and disorder is a typical origin of cluster spin glass behavior Mydosh_2015 . While further research is necessary to exactly identify the microscopic origin of relatively large ferromagnetic interactions in triangular lattice Ni-BIF, we can speculate on it. A conversion of weaker antiferromagnetic exchange interactions into one order of magnitude stronger ferromagnetic interactions has been achieved in a magnetic MOF when adding an electron on a linker molecule by doping. Perlepe2020 . The extra electron in the radical linker interacting antiferromagnetically with each of the nodes provides the FM exchange between the metal nodes. Along these lines, we can suggest a concentration of disordered radical spins on some of the linkers. Since each linker binds to three metal nodes in this structure, a single radical could create a cluster of ferromagnetically ordered spins that sits in a lattice of much weaker AFM interactions.

### v3

@@ Abstract
Magnetic metal-organic frameworks (MMOFs), where magnetic metal nodes are connected into a crystal structure by organic linkers, have a potential to host exotic magnetic states. We present a study of bulk magnetic properties of four metal-organic frameworks with the same boron imidazolate linkers, Cu-BIF, Co-BIF, Ni-BIF, and newly synthesized Zn-BIF, displaying a variety of lattice structures and nontrivial magnetic behaviors. While non-magnetic Zn-BIF provides an offset of magnetic response, magnetic susceptibility measurements of the other three magnetic materials demonstrate the presence of weak magnetic interactions in these MOFs, which differ between materials by sign and size. Cu-BIF, where magnetic nodes are connected into octahedral cages, shows simple paramagnetic behavior. Triangular lattice Co-BIF shows antiferromagnetic interactions on the order of 1 K, and a spin-crossover-like effect in magnetic susceptibility due to thermal depopulation of excited crystal electric field levels. Magnetic properties of Ni-BIF suggest sizable ferromagnetic interactions. Using DC/AC susceptibility and variable-field DC magnetization, we detect cluster spin-glass behavior in Ni-BIF and discuss possible microscopic origins of this behavior. This work demonstrates the variety of magnetic properties that are possible with a single organic ligand, and establishes the low energy scale of magnetic interactions through superexchange in boron imidazolate frameworks.
@@ I Introduction
Frustrated magnetism is one of the hot topics of current condensed matter research. While one of the ultimate goals is to synthesize and characterize spin liquid materials, where highly interacting spins are not ordered down to low temperatures Balents2010 ; Broholm2020 , many exotic magnetic states and new materials are found as a part of this research. One underexplored source of new magnetic materials are magnetic metal-organic frameworks (MMOFs), in which metal nodes are connected into crystals by organic linkers Thorarinsdottir2020 . Due to the versatility of organic chemistry and a relatively easy way to substitute metallic nodes without changing the structure or achieving different structures with the same magnetic nodes and linkers, these materials have the potential to be highly tunable magnets. There are a number of theoretical proposals on the realization of exotic magnetism in these materials Jacko2015 ; Yamada2017 ; yamada2017crystalline ; Jiang2021 ; zhang2017theoretical ; kambe2014redox , as well as a fair amount of MMOFs synthesized Thorarinsdottir2020 , but systematic studies of magnetic properties are rare.
In this manuscript, we present systematic magnetic studies of four boron imidazolate frameworks (BIFs), a unique subclass of MOFs that bear resemblances to zeolites and zeolitic imidazolate frameworks Zhang2016a ; Isaeva2020a , pictured in Fig. 1. Controlling the synthetic conditions as well as selecting the identity of the metal ion precursor can lead to a variety of solid-state structures. In this study, we selected 4 BIFs that are supported by the same boron trisimidazolate linker, but with different metal ions. Co-BIF and Ni-BIF are isostructural and form layered structures of 2D triangular lattices of MN6 octahedra (M = Co (J=1/2), Ni (S=1) ) with $D_{3d}$ symmetry connected by boron imidazolate molecules. Cu-BIF forms a M6L8 (M = metal ion, L = ligand) cage structure, with 6 Cu2+ ions (S=1/2) in an octahedral arrangement around a central pore. Finally, Zn-BIF forms a different cage structure (M4L4), with 4 Zn2+ (S=0) ions in a tetrahedral arrangement around the pore. These structures allow us to explore the effects of metal node substitution and MOF crystal structure on the magnetic properties of MMOFs.
@@ III.4 Ni-BIF
Figure 8: (a) In-phase component of AC susceptibility of Ni-BIF. $\chi^{\prime}$ data have been normalized to high-temperature $\chi^{\prime}$ at 999 Hz; raw data and scaling factors are presented in the SISI . (b) Vogel-Fulcher fit to $T_{f}$ obtained from $\chi^{\prime}$. (c) Power law fit to $T_{f}$ obtained from $\chi^{\prime}$.
To further understand the temperature-dependent spin relaxation dynamics of Ni-BIF we performed ZFC AC susceptibility measurements on the same sample. In-phase susceptibility $\chi^{\prime}$ is shown in Fig. 8a for 5 different frequencies with zero applied DC field. The $\chi^{\prime}$ data have been normalized to the high-temperature susceptibility at 999 Hz in order to clearly show the evolution of the freezing temperature $T_{f}$. Raw data are presented in Fig. 7b of the SI SI . $\chi^{\prime}$ displays a broad, complex peak shape consistent with a broad crossover in DC magnetic susceptibility (Fig. 5). The peak in $\chi^{\prime}$ shifts to higher temperatures as frequency is increased. The out-of-phase component $\chi^{\prime\prime}$, shown in Fig. 7c of the SI SI , highlights the complex peak shape. $T_{f}$ is approximated from the maximal point on the high-temperature side of the broad peak, as shown in Fig. 8a. $T_{f}$ changes from 11.7 K to 12.5 K between 1 Hz and 999 Hz; thus the Mydosh parameter, $\Delta T_{f}/(T_{f}\log_{10}(\Delta\nu))$, or the fractional change in $T_{f}$ per decade of frequency, is found to be 0.02. This parameter can be used to distinguish between spin-glass and superparamagnetic states, as in conventional spin-glasses it is found to be $\lesssim$ 0.01 while in superparamagnets it is $\gtrsim$ 0.1 Mydosh_2015 ; Bag2018 . An intermediate value such as that observed for Ni-BIF is attributed to cluster spin-glass behavior Bag2018 ; Banerjee2022_CSG .
An estimation of characteristic relaxation times allows to distinguish between spin glass and cluster spin glass Bag2018 ; Banerjee2022_CSG . Relaxation time can be estimated using the Vogel-Fulcher model (see Fig. 8b) which adapts the noninteracting Néel-Arrhenius law with a phenomenological parameter $T_{0}$ accounting for intercluster interactions Souletie1985 :
@@
where $\tau^{*}$ is a characteristic relaxation time similar to $\tau_{0}$, $T_{g}$ is the spin-glass temperature at zero frequency, and the power $zv^{\prime}$, which is treated as a single value for the fit, is the dynamical critical exponent for the relaxation time near the crossover. The fit in Fig. 8c suggests $\log_{10}(\tau^{*})=-6.8(7)$ ($\tau^{*}\sim 4.1(3.8)*10^{-7}$ s), $T_{g}$ = 11.62(4) K, and $zv^{\prime}$ = 2.8(6) K. The suggested ranges of the characteristic relaxation time and the dynamical critical exponent overlap with the ranges of $10^{-7}-10^{-10}$ s and 3-10 typically observed in cluster spin-glass systems Banerjee2022_CSG .
@@ IV Discussion
A combination of magnetic frustration, competing antiferro- and ferromagnetic interactions, and disorder is a typical origin of cluster spin glass behavior Mydosh_2015 . While further research is necessary to exactly identify the microscopic origin of relatively large ferromagnetic interactions in triangular lattice Ni-BIF, we can speculate on it. A conversion of weaker antiferromagnetic exchange interactions into one order of magnitude stronger ferromagnetic interactions has been achieved in a magnetic MOF when adding an electron on a linker molecule by doping. Perlepe2020 . The extra electron in the radical linker interacting antiferromagnetically with each of the nodes provides the FM exchange between the metal nodes. Along these lines, we can suggest a concentration of disordered radical spins on some of the linkers. Since each linker binds to three metal nodes in this structure, a single radical could create a cluster of ferromagnetically ordered spins that sits in a lattice of much weaker AFM interactions.